OpenStudy (anonymous):

Simplify by rationalizing the denominator. a/sqrt10+sqrt2

OpenStudy (mertsj):

Multiply the numerator and denominator by the conjugate of the denominator which is: \[\sqrt{10}-\sqrt{2}\]

OpenStudy (anonymous):

so this is the answer then 1/2 sqrt2/5a-sqrt2

OpenStudy (mertsj):

no

OpenStudy (anonymous):

then what did I do wrong

OpenStudy (mertsj):

\[\frac{a}{\sqrt{10}+\sqrt{2}}\times\frac{\sqrt{10}-\sqrt{2}}{\sqrt{10}-\sqrt{2}}=\frac{a \sqrt{10}-a \sqrt{2}}{10-2}\]

OpenStudy (anonymous):

oic thank you